OpenStudy (anonymous):

5 over 8 + Y=2 3 over 8 equations and raational numbers

OpenStudy (baldymcgee6):

lol.. theres a button called "equations", you should use it.

OpenStudy (baldymcgee6):

\[\frac{ 5 }{ 8 } + y = \frac{ 23 }{ 8 }\] I think this is what you mean

OpenStudy (anonymous):

yea but the 2 on the last one is a mixed number

OpenStudy (baldymcgee6):

oh, then that changes everything.. do you know what 2 3/8 =

OpenStudy (anonymous):

no

OpenStudy (anonymous):



OpenStudy (anonymous):

i just dont know how to solve it

OpenStudy (baldymcgee6):